a circle is inscribed in a square with sides 10cm long find the area of the shaded region in square centimiters.use 3.14 for pi.
Answer by venugopalramana(3286)

there is no diagram showing shaded area.assuming from the written part that it is the area between the square and circle
area of square=10*10=100
diameter of circle =10
radius=5
area=pi*R*r=3.14*5*5=78.5
so the difference is
100-78.5=22.5